The human ear canal is about 2.6 cm long. If it is regarded as a tube open at one end and closed at the eardrum, what is the fundamental frequency around which we would expect hearing to be most sensitive? 
Answer #1

fundamental frequency of one closed end pipe is given by:

f0 = n*V/(4*L)

V = Speed of sound = 343 m/s

L = length of ear canal = 2.6 cm = 0.026 m

n = 1, for fundamental frequency, So

f0 = 1*343/(4*0.026)

f0 = 3298.08 Hz = 3300 Hz

f0 = 3.3*10^3 Hz

f0 = 3.3 kHz

Let me know if you've any query.